OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

\[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x} and got -10x \sqrt{2x}\] what did she do wrong?

OpenStudy (compassionate):



OpenStudy (anonymous):

i just dont get this one problem :/

OpenStudy (dumbcow):

they didn't combine like terms correctly first simplify the radicals \[\sqrt{8x} = \sqrt{4}\sqrt{2x} = 2\sqrt{2x}\] \[\sqrt{18x} = \sqrt{9}\sqrt{2x} = 3\sqrt{2x}\] then you have \[3\sqrt{2x}+2x \sqrt{2x}-15\sqrt{2x} = 2x \sqrt{2x} -12\sqrt{2x}\] you can't combine them to make -10 because of the "x" outside the radical in one term